# A Comparative Study of Image Restoration Networks for General Backbone Network Design

### 24 Jan 2024 v2

@@ 4 General Backbone Network Design
Based on the above benchmark experiment, we believe that the principle of designing a general backbone network should be to ensure that the network can fulfill the functional requirements of all tasks. As Restormer shows relatively good task generality compared to the other networks, we select it as the starting point to design a more general network. By pinpointing and addressing the limitation of Restormer, we present an initial version of a general image restoration backbone network in this section.
Limitation of Restormer.
In the benchmark experiments, Restormer shows inferior performance to SwinIR on SR, particularly on Urban100. The qualitative comparison also indicates this phenomenon in Figure 4. Although Restormer can directly utilize global information, it fails to reconstruct fine textures even for self-repeated patterns, while SwinIR is capable of doing so. This discrepancy can be partially attributed to the U-shape architecture employed by Restormer. However, this architecture is integral to certain tasks. Another contributing factor is that Restormer relies on depth-wise convolution for spatial information interaction, and its spatial mapping capability is relatively weaker than spatial self-attention in SwinIR. To overcome this limitation, we incorporate a spatial self-attention module into Restormer to design a more powerful backbone network.
Network structure.
In Figure 5, we present our proposed backbone network, denoted as X-Restormer. We choose the U-shape architecture to build the network. In contrast to Restormer, we replace half of the transposed self-attention blocks (TSAB) with spatial self-attention blocks (SSAB) to enhance the ability of spatial information interaction. Given an input feature $F_{in}$, the two blocks process it alternately as:
@@
Table 3: Quantitative results on ×$\times 4$ image SR. * means the SwinIR model is pre-trained on ×$\times 2$ SR.
Table 4: Quantitative results on image denoising with the Gaussian noise level $\sigma=50$.
@@
Table 4: Quantitative results on image denoising with the Gaussian noise level $\sigma=50$.
Table 5: Quantitative results on motion deblurring.
@@
Specifically, we adopt the Multi-Dconv Transpose Attention (MDTA) as the TSA module. It first generates query (Q), key (K) and value (V) by applying $1\times 1$ convolutions followed by $3\times 3$ depth-wise convolutions. Then, the channel attention matrix of size $\mathbb{R}^{C\times C}$ is calculated by the dot-product of reshaped Q and K followed by a Softmax function. The schematic of TSA is shown in Figure 3. Finally, the result is generated by the dot-product of the attention matrix and V. For SSA, we adopt the Overlapping Cross-Attention (OCA) introduced in the HAT model [8]. We choose this approach because the shifted window mechanism is not suitable for our consecutive blocks, and HAT demonstrates the effectiveness of OCA. For the calculation of OCA, Q is produced by partitioning the input into non-overlapping windows, while K and V are generated by partitioning the input into overlapping windows with a manually set overlapping size. Apart from the different window partition methods, the calculation of OCA is essentially identical to that of standard window self-attention. For FFN, we employ the Gated-Dconv Feed-forward Network (GDFN) architecture, as used in Restormer. Instead of using two $1\times 1$ convolutions to construct an MLP, GDFN first processes input features through two $3\times 3$ depth-wise convolutions and $1\times 1$ convolutions. Then, the resulting features are combined via element-wise multiplication and pass through another $1\times 1$ convolution to produce the final output.
We have also tried multiple design choices for SSAB and TSAB, and the related experiments can be referred to Supp. We emphasize that our design of X-Restormer is not to develop novel architectures or modules to improve the performance on certain tasks, but to enhance the task generality of the network according to the principle of general backbone network design through as simple means as possible.
5 Experiments
5.1 Experimental Setup
We conduct experiments of the proposed X-Restormer on the same datasets used in the benchmark experiment. For the network implementation, the network employs a 4-level encoder-decoder with three times down-sampling and up-sampling. To maintain a similar number of parameters as Restormer, from level-1 to level-4, the numbers of consecutive blocks (containing a TSAB and a SSAB) are [2, 4, 4, 4], attention heads in TSA and SSA are both [1, 2, 4, 8], and channel numbers are [48, 96, 192, 384]. The channel expansion factor in GDFN is 2.66. The overall parameters are 26.06M, slightly less than Restormer of 26.13M. We adopt the same training settings as Restormer in the benchmark experiment to optimize the model. We use the AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.99$, utilizing an initial learning rate of $3e^{-4}$. The learning rate decay follows a cosine scheduler with intervals at 92k and 208k iterations, and the total training iterations are 300K. The input patch size is set to $256\times 256$ and the batch size is set to 32. For data augmentation, we use horizontal and vertical flips. We utilize the $L1$ loss function to train the model. Notably, we do not adopt any training tricks (e.g., ×$\times 2$ SR pretraining or EMA strategy) or testing tricks (e.g., TLC [11]).
@@ 5.1 Experimental Setup
×$\times 2$$\times 3$$\times 4$ $\sigma=15$ $\sigma=25$ $\sigma=50$
MPRNet 33.68/0.9301 27.92/0.8295 28.17/0.8043 34.27/0.9280 31.82/0.8901 28.60/0.8119 30.00/0.8812 31.20/0.9068 35.06/0.9874
SwinIR 33.83/0.9301 27.74/0.8351 28.14/0.8043 34.27/0.9283 31.83/0.8906 28.59/0.8143 29.06/0.8519 30.03/0.8983 31.48/0.9823
Uformer 29.99/0.8805 25.02/0.7622 27.88/0.7949 33.86/0.9254 31.42/0.8863 27.87/0.7891 29.64/0.8725 27.53/0.8569 29.92/0.9714
Restormer 34.51/0.9341 23.96/0.7410 28.70/0.8179 34.43/0.9303 32.02/0.8942 28.87/0.8222 30.54/0.8902 31.91/0.9134 36.95/0.9897
NAFNet 34.12/0.9314 23.28/0.7170 28.17/0.8087 34.18/0.9281 31.76/0.8908 28.64/0.8187 30.38/0.8911 31.56/0.9149 30.84/0.9797
X-Restormer 34.72/0.9360 28.04/0.8303 28.81/0.8217 34.67/0.9330 32.26/0.8983 29.12/0.8293 30.85/0.8983 32.27/0.9229 38.24/0.9914
5.2 Experimental Results
We compare our X-Restormer with the top three models in the benchmark experiments (based on the same test configurations) as well as other state-of-the-art approaches (based on the reported performance in their papers) in this section. PSNR/SSIM is provided in following tables. The best and second-best performance results are in bold and underline.
Image SR.
In Table 4, we present the quantitative results of ×$\times 4$ SR on five benchmark datasets: Set5 [2], Set14 [51], BSD100 [31], Urban100 [19] and Manga109 [33]. The state-of-the-art approaches, including RCAN [59], RCAN-it [27], SwinIR [35], IPT [4] and EDT [22] are compared in this experiment. X-Restormer significantly outperforms Restormer by 0.22dB on Set5, 0.34dB on Urban100 and 0.42dB on Manga109. This demonstrates the effectiveness of our design in enhancing the spatial mapping ability of Restormer. Furthermore, X-Restormer surpasses the SOTA method EDT by 0.2dB on Urban100 and 0.35dB on Manga109, indicating the effectiveness of X-Restormer on SR. Despite this, we point out that our method still cannot beat the most powerful SR approaches, such as HAT. We believe that in terms of the SR task, the plain residual in residual architecture designed for SR is still more effective.
Image denoising.
In Table 4, we provide the quantitative results of Gaussian denoising with the noise level $\sigma=50$ on four benchmark datasets: CBSD68 [32], Kodak24 [14], McMaster [58] and Urban100 [19]. The state-of-the-art methods: FFDNet [56], RNAN [61], RDN [62], IPT [4] and DRUNet [57] are compared in this experiment. As one can see that X-Restormer achieves the state-of-the-art performance, surpassing SwinIR by 0.36dB and outperforming Restormer by 0.22dB on Urban100. This demonstrates the superiority of X-Restormer on image denoising.
Image deblurring.
In Table 6, we compare the quantitative results of X-Restormer with the state-of-the-art methods: SPAIR [36], MIMO-UNet+ [10], IPT [4] and MPRNet [49] on both synthetic datasets (Gopro [34] and HIDE [39]) and real-world datasets (RealBlur-R and RealBlur-J [38]). X-Restormer achieves large performance gains over the other models on synthetic datasets, with an improvement of 0.36dB on Gopro compared to NAFNet and 0.54dB on HIDE compared to Restormer. Besides, our X-Restormer obtains the state-of-the-art performance on RealBlur-R and considerable performance on RealBlur-J, showing the effectiveness of our method on real-world motion deblurring scenarios.
Image deraining.
In Table 6, we present the quantitative results of deraining on benchmark datasets Test100 [18], Rain100L [47], Rain100H [47], Test1200 [52] and Test2800[16]. The state-of-the-art methods: PreNet[37], MSPFN [20], MPRNet [49] and SPAIR [36] are compared. X-Restormer outperforms the other models on Test100, Rain100H and Rain100L but performs inferior to Restormer on Test1200 and Test2800. This discrepancy is caused by the variations in degradation produced by different rain models. Nonetheless, our X-Restormer model exhibits comparable performance to state-of-the-art methods, demonstrating its effectiveness on image deraining.
Image dehazing.
In Table 7, we provide the quantitative results on the benchmark SOTS Indoor [21]. We compare the state-of-the-art approaches: PFDN [13], FFA-Net [46], AECR-Net [45], MAXIM [41] and DehazeFormer [40] in this experiment. Notably, our X-Restormer model significantly outperforms Restormer by a large margin of 0.93dB. When compared to the state-of-the-art dehazing method DehazeFormer, our X-Restormer achieves a breakthrough performance gain of 2.85 dB. These results demonstrate the superiority of our proposed method for image dehazing.
All-in-One Restoration.
@@ 5.2 Experimental Results
We conduct experiments on an all-in-one restoration setting, training different networks on five tasks with varying degradation levels (i.e., three SR upscaling factors and $\sigma\in(0,50)$ random Gaussian noises) simultaneously. The sampling probability for each task is the same and the average performance of benchmark datasets is calculated. As shown in Table 8, X-Restormer demonstrates significant advantages in handling multiple tasks concurrently, with its performance far exceeding other methods on all tasks. This indicates that a general backbone network is of great significance for building a general model that can handle multiple image restoration tasks, which can effectively mitigate task conflicts with the performance drops.
Summary.
@@
With enhanced spatial mapping capability, our X-Restormer achieves significant performance improvement compared to Restormer. Specifically, X-Restormer obtains large performance gains against Restormer of 0.42dB (Manga109), 0.22dB (Urban100), 0.54dB (HIDE), 0.61dB (Rain100H) and 0.93dB (SOTS Indoor) on image SR, denoising, deblurring, deraining and dehazing, showing the effectiveness of our design. Moreover, on all these five tasks, X-Restormer obtains state-of-the-art performance. Despite its simplicity, X-Restormer clearly has better task generality than other methods. Furthermore, we demonstrate that a powerful general backbone network has the potential to better handle multiple restoration tasks simultaneously. We hope it can inspire more works on the general image restoration backbone network design.

### 14 Mar 2024 v3

@@ 4 General Backbone Network Design
Based on the benchmark experiments, we believe that the principle of designing a general backbone network should be to ensure that the network can fulfill the functional requirements of all tasks. As Restormer shows relatively good task generality, we select it as the starting point to design a more general network. By pinpointing and addressing the limitation of Restormer, we present an initial version of a general image restoration backbone network in this section.
Limitation of Restormer. In the benchmark experiments, Restormer shows inferior performance to SwinIR on SR, particularly on Urban100. The qualitative comparisons also indicate this phenomenon in Fig. 4. From the visual and LAM [16] results, We can observe that Restormer can exploit large-range and even global information for the reconstruction. However, compared to SwinIR, it fails to reconstruct fine textures, even for self-repeated patterns, This discrepancy can be attributed to the U-shape architecture adopted by Restormer on the one hand, which increases the difficulty of reconstructing high-frequency information. On the other hand, Restormer relies on depth-wise convolution for spatial information interaction, whose spatial mapping capability is relatively weaker than the spatial self-attention in SwinIR. Considering that the U-shape architecture is indispensable for some tasks, we still need to retain this architectural design for task generality. To overcome the limitation of Restormer and design a more powerful backbone network, we choose to further enhance its spatial information interaction ability. An intuitive and feasible solution is to incorporate the spatial self-attention module into Restormer.
[…]
@@
Specifically, we adopt the Multi-Dconv Transpose Attention (MDTA) as the TSA module. It first generates query (Q), key (K) and value (V) by applying $1\times 1$ convolutions followed by $3\times 3$ depth-wise convolutions. Then, the channel attention matrix of size $\mathbb{R}^{C\times C}$ is calculated by the dot-product of reshaped Q and K followed by a Softmax function. The schematic of TSA is shown in Fig. 3. Finally, the result is generated by the dot-product of the attention matrix and V. For SSA, we adopt the Overlapping Cross-Attention (OCA) introduced in the HAT model [8]. We choose OCA because the shifted window mechanism in SwinIR is not intuitively suitable for our TSA-SSA consecutive blocks, and HAT demonstrates the effectiveness and superiority of OCA. For the specific calculation, Q is produced by partitioning the input into non-overlapping windows, while K and V are generated by partitioning the input into overlapping windows with a manually set overlapping size. Apart from the different window partition methods, the calculation of OCA is essentially identical to that of standard window self-attention. For FFN, we employ the Gated-Dconv Feed-forward Network (GDFN) architecture, as used in Restormer. Instead of using two $1\times 1$ convolutions to construct an MLP, GDFN first processes input features through two $3\times 3$ depth-wise convolutions and $1\times 1$ convolutions. Then, the resulting features are combined via element-wise multiplication and pass through another $1\times 1$ convolution to produce the final output.
We have also tried multiple design choices for SSAB and TSAB. Experiments can be found in Supp. We emphasize that our design of X-Restormer is not to develop novel architectures or modules to improve the performance on certain tasks, but to enhance the task generality of the network according to the principle of general backbone network design through as simple means as possible.
5 Experiments
5.1 Experimental Setup
We conduct experiments of the proposed X-Restormer on the same datasets used in the benchmark experiment. For the network implementation, the network employs a 4-level encoder-decoder with three times down-sampling and up-sampling. To maintain a similar number of parameters as Restormer, from level-1 to level-4 (i.e., $L_{1}\sim L_{4}$ in the figure) the numbers of consecutive blocks (containing a TSAB and a SSAB) are [2, 4, 4, 4] and the number of refinement blocks (i.e., $L_{r}$) is 4. Attention heads in TSA and SSA are both [1, 2, 4, 8], and channel numbers are [48, 96, 192, 384]. For OCA, the window size and the overlapping ratio are set to 8 and 0.5 as in HAT. The channel expansion factor in GDFN is 2.66. The overall parameters are 26.06M, slightly less than Restormer of 26.13M. We adopt the same training settings as Restormer in the benchmark experiment to optimize the model. We use the AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.99$, utilizing an initial learning rate of $3e^{-4}$. The learning rate decay follows a cosine scheduler with intervals at 92k and 208k iterations, and the total training iterations are 300K. The input patch size is $256\times 256$ and the batch size is 32. For data augmentation, we use horizontal and vertical flips. We utilize the $L1$ loss function to train the model. Notably, we do not adopt any training tricks (e.g., ×$\times 2$ SR pretraining or EMA strategy) or testing tricks (e.g., TLC [11]).
Table 3: Quantitative results on ×$\times 4$ image SR. * means the model pretrained on ×$\times 2$ SR.
Table 4: Quantitative results on image denoising with the noise level $\sigma=50$.
@@ 5.1 Experimental Setup
Table 4: Quantitative results on image denoising with the noise level $\sigma=50$.
Table 5: Quantitative results on image deblurring (motion blur).
@@
5.2 Experimental Results
We compare our X-Restormer with the top three models in the benchmark experiments (based on the same test configurations) as well as several state-of-the-art approaches for each task (based on the reported performance in their papers) in this section. PSNR(dB)/SSIM is provided in following tables. The best and second-best performance results are in bold and underline.
Image SR. In Tab. 4, we present the quantitative results of ×$\times 4$ SR on five benchmark datasets: Set5 [2], Set14 [50], BSD100 [31], Urban100 [18] and Manga109 [33]. The state-of-the-art approaches, including RCAN [57], RCAN-it [27], SwinIR [35], IPT [4] and EDT [22] are compared in this experiment. X-Restormer significantly outperforms Restormer by 0.22dB on Set5, 0.34dB on Urban100 and 0.42dB on Manga109. This demonstrates the effectiveness of our design in enhancing the spatial mapping ability of Restormer. Furthermore, X-Restormer surpasses the SOTA method EDT by 0.2dB on Urban100 and 0.35dB on Manga109, indicating the effectiveness of X-Restormer on SR. Despite this, we point out that our method still cannot beat the most powerful SR approaches, e.g., HAT. This is due to the inevitable weakening of SR performance for the U-shape architecture. In terms of SR, the plain residual in residual architecture is still more effective.
Image denoising. In Tab. 4, we provide the quantitative results of Gaussian denoising with the noise level $\sigma=50$ on four benchmark datasets: CBSD68 [32], Kodak24 [13], McMaster [56] and Urban100 [18]. The state-of-the-art methods: FFDNet [55], RNAN [58], RDN [60], IPT [4] and DRUNet [52] are compared in this experiment. X-Restormer achieves the state-of-the-art performance, surpassing SwinIR by 0.36dB and outperforming Restormer by 0.22dB on Urban100. This demonstrates the superiority of X-Restormer on image denoising.
Image deblurring. In Tab. 6, we compare the results of X-Restormer with the state-of-the-art methods: SPAIR [36], MIMO-UNet+ [10], IPT [4] and MPRNet [49] on both synthetic datasets (Gopro [34] and HIDE [39]) and real-world datasets (RealBlur-R and RealBlur-J [38]). X-Restormer achieves large performance gains over the other models on synthetic datasets, with an improvement of 0.36dB on Gopro compared to NAFNet [^5] By using TLC, on Gopro/HIDE, NAFNet: 33.69/31.32, X-Restormer: 33.89/31.87. and 0.54dB on HIDE compared to Restormer. Besides, our X-Restormer obtains the state-of-the-art performance on RealBlur-R and considerable performance on RealBlur-J, showing the effectiveness of our method on real-world motion deblurring scenarios.
Image deraining. In Tab. 6, we present the quantitative results of deraining on Test100 [19], Rain100L [47], Rain100H [47], Test1200 [51] and Test2800[15]. The state-of-the-art methods: PreNet[37], MSPFN [20], MPRNet [49] and SPAIR [36] are compared. X-Restormer outperforms the other models on Test100, Rain100H and Rain100L but performs inferior to Restormer on Test1200 and Test2800. This discrepancy is due to the variations in degradation produced by different rain models. Nonetheless, X-Restormer exhibits comparable performance to state-of-the-art methods, showing its effectiveness on image deraining.
Image dehazing. In Tab. 7, we provide the quantitative results on SOTS Indoor [21]. We compare the state-of-the-art approaches: PFDN [12], FFA-Net [46], AECR-Net [44], MAXIM [41] and DehazeFormer [40] in this experiment. Notably, X-Restormer model significantly outperforms Restormer by a large margin of 0.93dB. When compared to the state-of-the-art dehazing method DehazeFormer, our method achieves a breakthrough performance gain of 2.85 dB. These results demonstrate the superiority of X-Restormer for image dehazing.
@@ 5.2 Experimental Results
×$\times 2$$\times 4$ $\sigma=15$ $\sigma=25$ $\sigma=50$
MPRNet 33.68/0.9300 28.17/0.8043 34.27/0.9280 31.82/0.8901 28.60/0.8119 30.00/0.8812 31.20/0.9068 35.06/0.9874
SwinIR 33.83/0.9301 28.14/0.8043 34.27/0.9283 31.83/0.8906 28.59/0.8143 29.06/0.8519 30.03/0.8983 31.48/0.9823
Uformer 29.99/0.8805 27.88/0.7949 33.86/0.9254 31.42/0.8863 27.87/0.7891 29.64/0.8725 27.53/0.8569 29.92/0.9714
Restormer 34.51/0.9341 28.70/0.8179 34.43/0.9303 32.02/0.8942 28.87/0.8222 30.54/0.8902 31.91/0.9134 36.95/0.9897
NAFNet 34.12/0.9314 28.17/0.8087 34.18/0.9281 31.76/0.8908 28.64/0.8187 30.38/0.8911 31.56/0.9149 30.84/0.9797
X-Restormer 34.72/0.9360 28.81/0.8217 34.67/0.9330 32.26/0.8983 29.12/0.8293 30.85/0.8983 32.27/0.9229 38.24/0.9914
@@
All-in-One Restoration. We conduct experiments on an all-in-one restoration setting to show the effectiveness of different backbone networks in addressing various tasks simultaneously. Networks are trained on five tasks with varying degradation levels (i.e., ×$\times 2$, ×$\times 4$ upscaling factors for SR and $\sigma\in(0,50)$ random noise level for Gaussian denoising). The sampling probability for each kind of task is the same, and the average performance of common benchmark datasets (appearing in Tab. 4 $\sim$ Tab. 7) is calculated. As shown in Tab. 8, with the relatively better task generality among the existing networks, Restormer exhibits better performance on the all-in-one restoration. By overcoming the limitation of Restormer, our X-Restormer demonstrates further advantages in handling multiple tasks concurrently, with its performance far exceeding other networks on all tasks. In contrast, the other networks are more or less affected by optimization conflicts across different tasks (e.g., SwinIR performs inferior to Restormer even on SR). These indicate that a general backbone network is of great significance for building a general model that process multiple image restoration tasks, which can effectively mitigate task conflicts with the performance drops.
@@
Summary. With enhanced spatial mapping capability, our X-Restormer achieves significant performance improvement compared to Restormer. Specifically, X-Restormer obtains performance gains against Restormer of 0.42dB (Manga109), 0.22dB (Urban100), 0.54dB (HIDE), 0.61dB (Rain100H) and 0.93dB (SOTS Indoor) on image SR, denoising, deblurring, deraining and dehazing, respectively, showing the effectiveness of our design. Moreover, on all these five tasks, X-Restormer obtains state-of-the-art performance. Despite its simplicity, X-Restormer clearly has better task generality than other methods. Furthermore, we demonstrate that a powerful general backbone network has the potential to better handle multiple restoration tasks simultaneously. We hope it can inspire more works on the general image restoration backbone network design.